OpenStudy (anonymous):

Perform the indicated operation. Leave all answers in the form a + b i. (7 - 2 i ) - (6 - 3 i ) a. 1-i b. 1+5i c. 1+i d. 1-5i

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

think that \[(7 - 2 i ) - (6 - 3 i ) \] is just like \[(7 - 2 x) - (6 - 3 x ) \]

Parth (parthkohli):

Treat \(i\) like a variable. These imaginary do not deserve special treatment. lol

OpenStudy (anonymous):

It's not like they get a whole field in mathematics to themselves...

OpenStudy (anonymous):

clear what to do right? remove parentheses, don't forget the distributive law, then combine like terms

OpenStudy (unklerhaukus):

\[(7 - 2 i ) - (6 - 3 i ) \]\[=7 - 2 i -6 + 3 i \]\[=(7 -6)+ (- 2 + 3 )i\]

OpenStudy (anonymous):

c.?

OpenStudy (anonymous):

Or is it a.?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

The answer I got is 2i but that doesn't correlate to the options given

OpenStudy (unklerhaukus):

\[=(7−6)+(−2+3)i\]what does the first bracket simplify to ?

OpenStudy (unklerhaukus):

7-6=1 right?

OpenStudy (unklerhaukus):

\[=1+(−2+3)i\]what about the second bracket? -2+3=?